I need help verifying this identity.
cot^2x-cos^2x = (cot^2x)(cos^2x)
So I tried:
(cos^2x)/(sin^2x) -cos^2 = (cos^2x)/(sin^2x)(cos^2x)
But after that I'm kind of stumped. Any help with future steps would be appreciated!
Thanks so much!

Answer by stanbon(75887)

cot^2x-cos^2x = (cot^2x)(cos^2x)
So I tried:
[(cos^2x)/(sin^2x)] - cos^2 = [(cos^2x)/(sin^2x)](cos^2x)
---------
Multiply both sides by sin^2(x) to get:
cos^2(x) - sin^2(x)*cos^2(x) = cos^4(x)
(cos^2(x)(1 - sin^2(x)) = cos^4(x)
(cos^2(x))(cos^2(x)) = cos^2(x)*cos^2(x)
-----------------------
Cheers,
Stan H.